OpenStudy (anonymous):

Find the third term of each arithmetic sequence: 1. A(n) = 8 + (n-1)(-3) 2. A(n) = -3(n-2) + 4

OpenStudy (anonymous):

well if u simplify 1 u get -3n+11 if u simplify 2 u get -3n+10 so next term shld be -3n+9, take -3 common u get -3(n-3) so ans is 3. A(n) = -3(n-3)

OpenStudy (anonymous):

silly me

OpenStudy (anonymous):

\[A(n)=8+(n-1)\times -3=-3n+11\] if this is written correctly

OpenStudy (anonymous):

so \[A(1)=-3\times 1+11=8\] \[A(2)=-3\times 2+8=2\]\[A(3)=-3\times 3+8=-1\]